# On the Effectiveness of Microservices Tactics and Patterns to Reduce Energy Consumption: An Experimental Study on Trade-Offs

### v1

@@ IV-A Effectiveness of Reducing Energy Consumption (RQ1)
To confirm if these mean differences were statistically significant, we performed t-tests and calculated Cohen’s $d$ for significant differences to estimate the effect size. We show the testing results, the effect sizes, and the mean percentage decrease in energy consumption in Table II. As suspected from the visualization, nearly no technique significantly impacted energy consumption at low load. The only exception was Use different proxies (T3), which had a small effect (Cohen’s $d$ of $-0.407$). Distribute pods to different nodes (T1) significantly reduced energy consumption for both medium load (large effect) and high load (moderate effect), indicating that load balancing can lead to a more energy-efficient system. Distribute microservices (T2) had a moderate effect on energy consumption for both medium and high load, with simplifying and reusing infrastructure contributing to efficiency. In addition to its small effect for low load, Use different proxies (T3) also significantly reduced energy consumption under medium load (large effect) and high load (moderate effect), suggesting that proxies play a vital role as load increases. Backends for Frontends (P1) showed a significant reduction for the higher load levels, with a large effect for medium and a moderate one for high loads, due to the benefits of fewer requests. Request Bundle (P2) showed a moderate effect for both medium and high loads, as the overhead of establishing and closing communications was reduced. Finally, Caching of Read Requests (P3) also significantly reduced energy consumption with a moderate effect for both medium and high loads, as fewer microservices needed to query databases and other services.
@@ IV-B1 Response Time
medium $<$0.0001 -3.989 -59.0%
high $<$0.0001 -4.955 -52.3%
P3: Caching of Read Requests low $<$0.0001 -4.046 -52.0%
medium $<$0.0001 -5.267 -81.5%
high $<$0.0001 -7.338 -72.5%
@@ IV-B2 Throughput
Lastly, Request Bundle (P2) performed most poorly for throughput on paper, with large negative effects at all load levels. We defined throughput as the number of requests handled by the system within a certain time span, which is a standard definition used by many studies [42, 43] as well as Locust. Therefore, this reduction is caused by grouping multiple messages together, which naturally reduced the number of served requests in the measured timeframe. In our implementation, Request Bundle aggregated the messages of three operations in every request and response. As a consequence, Request Bundle would the best choice among all techniques if we changed the definition of throughput to the number of handled operations within a timeframe (three times the baseline).
@@
P2: Request Bundle low $<$0.0001 -24.038 -60.5%
medium $<$0.0001 -39.996 -55.3%
high $<$0.0001 -19.423 -44.9%
@@ IV-C Impact on Maintainability (RQ2b)
TABLE VI: Summarized Increase (from $+$ to +++++++ + +) or Decrease (from -- to ---- - -) in Dependent Variables Through Tactics and Patterns
@@
Medium --- - n/a n/a worse
High ---- - - +++++ + --- - worse
T3: Use different proxies Low -- ---- - - n/a n/a
Medium ---- - - --- - +++++++ + + n/a
High ---- - - --- - +++++++ + + n/a
@@
Medium ---- - - n/a $+$ worse
High --- - $+$ n/a worse
P2: Request Bundle Low n/a +++++ + ---- - - better
Medium --- - ---- - - ---- - - better
High ---- - - ---- - - ---- - - better
P3: Caching of Read Requests Low n/a ---- - - $+$ worse
Medium --- - ---- - - +++++++ + + worse
High ---- - - ---- - - +++++++ + + worse
@@ V Discussion
In this section, we discuss important findings and the implications of our results. Table VI provides an aggregated summary of how strongly the tactics and patterns impacted our studied dependent variables, which can serve as a quick overview to identify the most useful techniques and potential trade-offs. Note that -- and $+$ mean the decrease and increase in value, not the nature of the impact. For example, ---- - - for the energy consumption or response time impact means a large decrease, which is beneficial. Cells with n/a indicate no difference between the treatment and the control group.
@@ VI Threats to Validity
We discuss potential threats to the validity of our results based on the dimensions described by Wohlin et al. [22]. Regarding construct validity, one typical threat for controlled experiments is potentially unreliable or non-representative measurements of the dependent variables. For example, we used Scaphandre to measure power at a specific time interval to derive energy consumption from it. We did so once every second to have a decently small interval for accurate energy consumption measurements. However, it should be noted that the result cannot be treated as continuous and smooth data. Nonetheless, it is a common practice to approximate the energy consumption in this way, e.g., as is the case for the Running Average Power Limit (RAPL) algorithm that is widely adopted by Intel [44]. We also tested Scaphandre extensively and observed that the power for the same treatment was very stable and reliable. We did not use smaller time intervals (less than one second) because it might have led to a heavier burden on the experiment server.
Another threat is related to our maintainability measurements. While the used maintainability sub-variables are based on ISO 25010 and Sigrid has been widely evaluated in industry, our source-code-focused measurements did not reflect the whole scope of maintainability, especially given the architecture-centric nature of microservices. As discussed in the results, the selected maintainability sub-variables could not detect some essential changes in our system, like the proxy configuration or node configuration change. While reduced maintenance efforts are generally seen as a benefit of Istio [26], this was not reflected in the source code measurements.
@@
Lastly, the single system we experimented with, Online Boutique, is an open-source e-commerce application, and naturally does not represent all types of microservice-based systems. For example, it contains around 5k LOC distributed over 9 services plus one Web frontend, making it fairly small compared to many industry applications [23]. Despite many similarities, open-source projects cannot fully reflect all characteristics of closed-source industry software. While we still believe in decent generalizability of our results, it is important to reproduce them across different systems and especially in real-world industry case studies.

### v2

@@ IV-A Effectiveness of Reducing Energy Consumption (RQ1)
To confirm if these mean differences were statistically significant, we performed t-tests and calculated Cohen’s $d$ for significant differences to estimate the effect size. We show the testing results, the effect sizes, and the mean percentage decrease in energy consumption in Table II. As suspected from the visualization, nearly no technique significantly impacted energy consumption at low load. The only exception was Use different proxies (T3), which had a small effect (Cohen’s $d$ of 0.407-0.407). Distribute pods to different nodes (T1) significantly reduced energy consumption for both medium load (large effect) and high load (moderate effect), indicating that load balancing can lead to a more energy-efficient system. Distribute microservices (T2) had a moderate effect on energy consumption for both medium and high load, with simplifying and reusing infrastructure contributing to efficiency. In addition to its small effect for low load, Use different proxies (T3) also significantly reduced energy consumption under medium load (large effect) and high load (moderate effect), suggesting that proxies play a vital role as load increases. Backends for Frontends (P1) showed a significant reduction for the higher load levels, with a large effect for medium and a moderate one for high loads, due to the benefits of fewer requests. Request Bundle (P2) showed a moderate effect for both medium and high loads, as the overhead of establishing and closing communications was reduced. Finally, Caching of Read Requests (P3) also significantly reduced energy consumption with a moderate effect for both medium and high loads, as fewer microservices needed to query databases and other services.
@@ IV-B1 Response Time
medium <<0.0001 -3.989 -59.0%
high <<0.0001 -4.955 -52.3%
P3: Caching of Read Requests low <<0.0001 -4.046 -52.0%
medium <<0.0001 -5.267 -81.5%
high <<0.0001 -7.338 -72.5%
@@ IV-B2 Throughput
Lastly, Request Bundle (P2) performed most poorly for throughput on paper, with large negative effects at all load levels. We defined throughput as the number of requests handled by the system within a certain time span, which is a standard definition used by many studies [de2016architecture, menasce2002load] as well as Locust. Therefore, this reduction is caused by grouping multiple messages together, which naturally reduced the number of served requests in the measured timeframe. In our implementation, Request Bundle aggregated the messages of three operations in every request and response. As a consequence, Request Bundle would the best choice among all techniques if we changed the definition of throughput to the number of handled operations within a timeframe (three times the baseline).
@@
P2: Request Bundle low <<0.0001 -24.038 -60.5%
medium <<0.0001 -39.996 -55.3%
high <<0.0001 -19.423 -44.9%
@@ IV-C Impact on Maintainability (RQ2b)
TABLE VI: Summarized Increase (from ++ to ++++++) or Decrease (from - to ---) in Dependent Variables Through Tactics and Patterns
@@
Medium -- n/a n/a worse
High --- ++++ -- worse
T3: Use different proxies Low - --- n/a n/a
Medium --- -- ++++++ n/a
High --- -- ++++++ n/a
@@
Medium --- n/a ++ worse
High -- ++ n/a worse
P2: Request Bundle Low n/a ++++ --- better
Medium -- --- --- better
High --- --- --- better
P3: Caching of Read Requests Low n/a --- ++ worse
Medium -- --- ++++++ worse
High --- --- ++++++ worse
@@ V Discussion
In this section, we discuss important findings and the implications of our results. Table VI provides an aggregated summary of how strongly the tactics and patterns impacted our studied dependent variables, which can serve as a quick overview to identify the most useful techniques and potential trade-offs. Note that - and ++ mean the decrease and increase in value, not the nature of the impact. For example, --- for the energy consumption or response time impact means a large decrease, which is beneficial. Cells with n/a indicate no difference between the treatment and the control group.
@@ VI Threats to Validity
We discuss potential threats to the validity of our results based on the dimensions described by Wohlin2024. Regarding construct validity, one typical threat for controlled experiments is potentially unreliable or non-representative measurements of the dependent variables. For example, we used Scaphandre to measure power at a specific time interval to derive energy consumption from it. We did so once every second to have a decently small interval for accurate energy consumption measurements. However, it should be noted that the result cannot be treated as continuous and smooth data. Nonetheless, it is a common practice to approximate the energy consumption in this way, e.g., as is the case for the Running Average Power Limit (RAPL) algorithm that is widely adopted by Intel [hahnel2012measuring]. We also tested Scaphandre extensively and observed that the power for the same treatment was very stable and reliable. We did not use smaller time intervals (less than one second) because it might have led to a heavier burden on the experiment server.
Another threat is related to our maintainability measurements. While the used maintainability sub-variables are based on ISO 25010 and Sigrid has been widely evaluated in industry, our source-code-focused measurements did not reflect the whole scope of maintainability, especially given the architecture-centric nature of microservices. As discussed in the results, the selected maintainability sub-variables could not detect some essential changes in our system, like the proxy configuration or node configuration change. While reduced maintenance efforts are generally seen as a benefit of Istio [IstioImpact], this was not reflected in the source code measurements.
@@
Lastly, the single system we experimented with, Online Boutique, is an open-source e-commerce application, and naturally does not represent all types of microservice-based systems. For example, it contains around 5k LOC distributed over 9 services plus one Web frontend, making it fairly small compared to many industry applications [Bogner2019]. Despite many similarities, open-source projects cannot fully reflect all characteristics of closed-source industry software. While we still believe in decent generalizability of our results, it is important to reproduce them across different systems and especially in real-world industry case studies.